OpenStudy (anonymous):

a puzzle we have to place 30 birds on 5 trees such that no tree is left vacant and number of birds on each tree is odd ??? our maths teacher asked this , ....

OpenStudy (jamesj):

There is no solution to this problem. To see this, let n be an odd number. Then there is another integer k such that n = 2k + 1. For example 5 = 2 x 2 + 1. So suppose the five trees have n1, n2, n3, n4 and n5 birds respectively. Then write n1 = 2k1 + 1, n2 = 2k2 + 1, etc. Then the total n1 + n2 + n3 + n4 + n5 = 30 But also n1 + n2 + n3 + n4 + n5 = 2(k1 + k2 + k3 + k4 + k5) + 5 = 2(k1 + k2 + k3 + k4 + k5 + 2) + 1 i.e., the sum is an odd number. But 30 is not odd. Contradiction! Hence it is not possible that each tree have an odd number of birds.

OpenStudy (turingtest):

@JAmesJ Is that a proof by contradiction and induction, or just contradiction?

OpenStudy (jamesj):

Proof by contradiction: we want to prove a statement A. Assume A is false and show that leads to a logical contradiction. Then it must be the case that A is true. In this case, A is the statement: "It is NOT possible to place 30 birds on 5 trees such that no tree is left vacant and number of birds on each tree is odd."

OpenStudy (turingtest):

I saw it was at least by contradiction, but by representing odds as 2k+1... oh, I just realized I thought somehow you slipped a k+1 in there which is what made me think of induction. Never mind, thanks for clearing that up.